Wednesday, April 17, 2013

Night Log 2013-04-16

SA: Tim
SO: Thea
Others: alexei


entire night lost to clouds, wind, and then humidity as well. spent the evening analyzing FP calibration data and then testing the new numbers out. MR is working well with the new calibration in place, though some bugs were encountered in PCON's FP calibration routine. also checked the MR calibration near 5000 A and it appears to not have changed much. the lines i tried appeared where they were supposed to within a few A. calibration parameters for TF were also updated and extended out to 7500 A.

No comments:

Post a Comment